# Ladder-residual: parallelism-aware architecture for accelerating large model inference with communication overlapping

### v2

@@ Abstract
Large language model inference is both memory-intensive and time-consuming, often requiring distributed algorithms to efficiently scale. Various model parallelism strategies are used in multi-gpu training and inference to partition computation across multiple devices, reducing memory load and computation time. However, using model parallelism necessitates communication of information between GPUs, which has been a major bottleneck and limits the gains obtained by scaling up the number of devices. We introduce Ladder Residual, a simple architectural modification applicable to all residual-based models that enables straightforward overlapping that effectively hides the latency of communication. Our insight is that in addition to systems optimization, one can also redesign the model architecture to decouple communication from computation.. While Ladder Residual can allow communication-computation decoupling in conventional parallelism patterns, we focus on Tensor Parallelism in this paper, which is particularly bottlenecked by its heavy communication. For a Transformer model with 70B parameters, applying Ladder Residual to all its layers can achieve 30% end-to-end wall clock speed up at inference time with TP sharding over 8 devices. We refer the resulting Transformer model as the Ladder Transformer. We train a 1B and 3B Ladder Transformer from scratch and observe comparable performance to a standard dense transformer baseline. We also show that it is possible to convert parts of the Llama-3.1 8B model to our Ladder Residual architecture with minimal accuracy degradation by only retraining for 3B tokens. To further push the performance frontier, we propose another architectural modification that directly drops communications in the model, unlocking fast LLM inference in settings devoid of NVLink or other fast interconnects. We release our code for training  [^1] Training code: https://github.com/IBM/dolomite-engine and inference  [^2] Inference code: https://github.com/mayank31398/ladder-residual-inference for easier replication of experiments.
**footnotetext: Equal contribution, correspondence to muruzhan@usc.edu, mayank.mishra2@ibm.com, tri@tridao.me
@@ 1 Introduction
With the rapid scaling of Large Language Models (LLMs) (Smith et al., 2022; Workshop et al., 2023; Brown, 2020), the compute and memory requirements for training and inference have grown significantly. Tensor parallelism (TP) (Shoeybi et al., 2020) is a widely used model parallelism technique that partitions the weights and intermediate activations across multiple GPUs. In contrast to pipeline parallelism (Narayanan et al., 2021) and data parallelism (Li et al., 2020), which rely on processing independent batches of user requests on each device, tensor-parallel inference enables multiple devices to cooperate to process a single batch of user requests at a time, therefore in theory allowing infinite scaling given a sufficient number of processors. However, TP requires synchronizing the partitioned intermediate activations across the GPUs. This synchronization is a blocking AllReduce operation on the activations across the GPUs and is therefore bottlenecked by the network communication latency. Even for GPUs connected via fast interconnects (like NVLink (NVIDIA Corporation, 2024)), the communication costs can account for 38% of the latency at inference time when running a 70B transformer with batch size 4 and TP world size of 8.
@@
Communication is blocking because there is a sequential structure between communication and computation in existing model designs: we wait for communication in order to prepare the correct input for the next computation. In the prevalent residual-based architectures, the computation flow can be written as $x_{i+1}=h_{i+1}(x_{i})+x_{i}$, where $x_{i}$ is the residual stream after layer $i$ and $h_{i+1}$ is the computation at layer $i+1$. Notice that the communication of $x_{i}$ needs to be done before executing $h_{i+1}$ if $h_{i}$ is partitioned across devices. Liu et al. (2023b) found that activation changes slowly in Transformer, as the norm of each update $h_{i+1}(x_{i})$ is small compared to the residual. Based on this observation, we hypothesize that maintaining the regular residual connection is enough to restrict the representation shift, and we can feed each module a “stale” input to create overlapping opportunities.
We propose Ladder Residual, a simple change where we reroute the residual stream after module $i-1$ (instead of module $i$) as input to module $i+1$: $x_{i+1}=h_{i+1}(x_{i-1})+x_{i}$. With this design, the computation of $h_{i+1}$ is decoupled from the communication of $x_{i}$, enabling straightforward overlapping to hide the latency of communication. Figure 1 shows how Ladder Residual can be applied to the Transformer architecture. At inference time with TP world size of 8 (i.e., across 8 devices), running a 70B Transformer with Ladder Residual can be around 30% faster than the standard Transformer. In Table 1, we provide the inference speedup on Transformers of different sizes. The proposed Ladder Residual method can also be used to accelerate other forms of parallelism, although we focus on Tensor Parallelism in this paper for inference as it is particularly bottlenecked by the heavy communication. Our method obtains 5-7% training speedup when training an 8B model with 8k context length on 64 H100s with 3D parallelism across the GPUs (Tensor Parallel, Sequence Parallel and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023; Rajbhandari et al., 2020)), but we decide to focus on the inference speed ups in this paper since training with pure FSDP is usually faster because weights can be pre-fetched and gradient synchronization using ReduceScatter can be overlapped in FSDP making communications in FSDP non-blocking.
Table 1: Inference speedup from applying Ladder Residual on a Transformer model. The test setup is 1024 prompt length, 512 generated tokens, batch size 4, and TP world size of 8 for 1B, 3B, 8B, 34B, 70B, 176B model and TP world size of 16 for 405B model (using two nodes). The speedup value is calculated by comparing Ladder Transformer with Standard Transformer’s inference throughput in tokens per second. We measure the speedup both with and without NVLink interconnect.
Model size No NVLink With NVLink
@@
Because of the widespread use of Transformer (Vaswani, 2017) based Language models, we focus on applying Ladder Residual on Transformer models in this paper, and we call the resulting model as Ladder Transformer. However, it should be noted that Ladder Residual is compatible with any residual based model architecture. We conduct experiments under two scenarios to verify if we can maintain the same performance after applying Ladder Residual to a Transformer-based model:
@@
  • Post-training adaptation: We take the pretrained Llama-3.1-8B-instruct model (Dubey et al., 2024) and apply Ladder Residual on the upper half layers. We then fine-tune it on 3B tokens to adapt to the representation shift. With this relatively light retraining, we can obtain a hybrid Ladder Llama that is on par with the original Llama model on AlpacaEval, math, commonsense reasoning, and knowledge-intensive tasks while only being slightly worse on coding benchmarks.
In section 5, as a proof of concept of what Ladder Residual can unlock, we additionally propose Desynced Residual, an architectural modification that eliminates communication by restricting certain modules to act on the local activations on each device independently. We found that it’s possible to eliminate 75% of the communication without much performance degradation when training a 1B model. Since communication can be too heavy to overlap in some scenarios (for example, with cross-node communication or even wireless communication), applying Ladder Residual can further unlock the potential of scaling TP to a giant GPU cluster.
In our paper, we propose to change the model architecture to accelerate model parallelism without touching low-level kernels, making it easily deployable on any hardware. We show that such architecture modification performs on-par with the standard transformer. As model size grows, multi-gpu or even cross-node serving will become more and more important, and our research provides a fresh perspective on designing a model architecture with parallelism optimization in mind. Such design can be applied to any architecture that suffers from blocking communication, although in this paper we conduct experiments on Transformer-based language models due to their widespread use and popularity.
2 Background
In this section, we provide some background of Tensor Parallelism (TP) (Shoeybi et al., 2020) for a transformer-based language model.
2.1 Tensor parallelism
Tensor parallelism (Shoeybi et al., 2020) is a commonly used technique in distributed training/inference. It partitions weights and activations across devices and performs partial computations on each device. Consider a sequence of 2 linear layers with weight matrices $A$ and $B$ and input activation $X$ that is running on 2 GPUs (TP world size of 2), we split $A$ along the output dimension into $[A_{1},A_{2}]$, and split $B$ along the input dimension into $\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}$. Then the output of the sequence of the 2 linear layers can be computed as $(XA)B=(XA_{1})B_{1}+(XA_{2})B_{2}$ and we effectively partition the computation on the two devices. The final summation requires an AllReduce operation to aggregate the partial sums on each device, which introduces communication overhead. The AllReduce overhead increases with increasing message size and increasing number of devices participating in the AllReduce. A transformer layer consists of an attention block and an MLP block: both can be considered as a sequence of two matrix multiplications and therefore fit into the tensor parallelism paradigm described above. Thus each transformer layer contains 2 AllReduce operations: one for attention and another for MLP. Denoting the input to the $i^{th}$ block as $x_{i-1}$, the transformer can be viewed as the following sequential structure:
@@ 2.1 Tensor parallelism
where the * denotes a partial-sum that requires an AllReduce for the full output replicated across all the GPUs. Note that the AllReduce operation is the identity function for a model running on 1 GPU.
A Transformer with N layers needs to perform the AllReduce 2N times and this can account for 30% of the inference latency for a 70B model using TP world size of 8, even with NVLink interconnect. For communication without NVLink support or inter-node communication, their latency can account for over 50% of the end-to-end latency. Modern nodes contain GPUs connected via NVLink but usually have no more than 8 GPUs per node, due to limited PCIe lanes or power and cooling constraints in datacenters. There is a steep falloff in communication bandwidth and sharp increase in latency when the communication happens across nodes either over InfiniBand or Ethernet thus making scaling TP practically infeasible outside a node.
3 Ladder Transformer
In this section, we introduce the Ladder Residual architecture when applied to a Transformer model and benchmark its efficiency gain under various model sizes and generation setups.
Algorithm 1 Ladder Transformer Layer with Tensor Parallelism. Note that the AsyncAllReduce returns a handle which is passed to the next layer.
1:function Layer(residual,  prev_attn_out,  prev_mlp_out,                prev_attn_handle,  prev_mlp_handle)
2:     prev_attn_handle.wait()
@@ 3 Ladder Transformer
13:
14:     return residual,  attn_out,  mlp_out,  attn_handle,  mlp_handle
15:end function
@@ 3.1 Architecture description
Note that the residual stream of each block still takes the output from the previous module as usual, this ensures the module $i$ is still able to process information from all previous $i-2$ modules.
@@
NVLINK-UpperBound-Llama-70B 30.54 30.00 42.90
NVLINK-Parallel-Llama-70B 5.42 18.04 21.75
NVLINK-Ladder-Llama-70B 5.78 23.71 30.79
NO-NVLINK-UpperBound-Llama-70B 35.84 52.71 110.7
NO-NVLINK-Parallel-Llama-70B 14.92 28.73 40.07
NO-NVLINK-Ladder-Llama-70B 6.94 37.71 59.87
@@ 3.2 Inference Implementation
Alignment with Real-World Scenarios: To evaluate the practical benefits of Ladder Residual, we integrated this mechanism into a standard Llama-like Transformer. Building upon gpt-fast (PyTorch Labs, 2024), we partition the weights of the attention and feedforward modules for tensor parallelism to optimize inference speed. We use CUDA graphs (Coleman, 2020) via PyTorch compile (with “reduce-overhead” mode) to generate static computation graphs for both the prefill and decode phase to reduce CPU kernel launch overheads which can be a big bottleneck especially during the decode phase of Transformer inference. PyTorch compile also additionally helps in accelerating inference and reducing the memory consumption of the models by emitting more efficient kernels.
Figure 2: Improvement in end-to-end inference throughput achieved by communication-efficient architectures relative to a traditional transformer, benchmarked on Llama-3 70B. Standard refers to the regular Llama-3, and Ladder is Llama-3 with our Ladder Residual architecture. Running with NVLink, Ladder Residual architecture can achieve up to $29\%$ greater throughput than the traditional Transformer. Without NVLink, we observe speedups up to $60\%$. All experiments were conducted on a generation task with $1024$ prompt tokens and $512$ completion tokens. Missing data points indicate CUDA OOM.
@@ 3.3 Faster Inference with Ladder Residual
In this section, we benchmark Ladder Residual under various scenarios and show that across various model sizes, batch sizes, and TP world sizes, Ladder Residual can obtain considerable speedup over the standard Transformer. We also benchmark under the case where NVLink is disabled, and show that our method can obtain more than 50% of the speedup when high-speed interconnect is not available. Finally, we consider cross-node Tensor Parallelism, which can be necessary for serving large models like Llama 3.1 405B, and show that Ladder Residual can improve inference throughput by more than $30\%$.
@@ 3.3.1 Setup
  • Parallel Attention and MLP: Following the PaLM parallelization strategy Chowdhery et al. (2022); Wang & Komatsuzaki (2021), we fuse the weights of the query, key, value, gate, and up projections into a single matrix. The outputs are then split, and the attention and swiglu are performed in parallel. While not proposed for accelerating Tensor Parallelism, we realize this architecture can effectively cut half of the communication, with the extra benefit of being able to fuse attention and mlp together, therefore we consider it as an alternative to Ladder Residual.
@@
Note that different Transformer-based models usually have slightly different designs. We choose to use Llama-3’s model architecture as a base since its one of the most widely used models. The benchmarking results here are a good representative of a variety of different design choices since the communication patterns are mostly the same across various transformer variants.
To simulate various inference scenarios, we select multiple experimental configurations. The prompt length and generation length is fixed to 1024 and 512 respectively, while we vary the tensor parallel world sizes among 1, 2, 4, and 8, and batch sizes among 1, 4, 16, and 64 to understand performance under different generation settings. All benchmarks are done on NVIDIA H100 GPUs.
To evaluate the impact of hardware communication capabilities, we adjust the NVLink settings using environment variables. We enable NVIDIA SHARP (Graham et al., 2016) by setting NCCL_NVLS_ENABLE=1 and disable NVLINK communication by setting NCCL_P2P_DISABLE=1. This allows us to assess the performance of different algorithms in varying communication environments.

### v4

@@ Abstract
Large language model inference is both memory-intensive and time-consuming, often requiring distributed algorithms to efficiently scale. Various model parallelism strategies are used in distributed training and inference to partition computation across multiple devices, reducing memory load and computation time. However, using model parallelism necessitates communication of information between GPUs, which has been a major bottleneck and limits the gains obtained by scaling up the number of devices. We introduce Ladder Residual, a simple architectural modification applicable to all residual-based models that enables straightforward overlapping that effectively hides the latency of communication. Our insight is that in addition to system optimizations, one can also redesign the model architecture to decouple communication from computation. While Ladder Residual can allow communication-computation decoupling in conventional parallelism patterns, we focus on Tensor Parallelism in this paper, which is particularly bottlenecked by its heavy communication. For a Transformer model with 70B parameters, applying Ladder Residual to all its layers can achieve 29% end-to-end wall clock speed up at inference time with TP sharding over 8 devices. We refer the resulting Transformer model as the Ladder Transformer. We train a 1B and 3B Ladder Transformer from scratch and observe comparable performance to a standard dense transformer baseline. We also show that it is possible to convert parts of the Llama-3.1 8B model to our Ladder Residual architecture with minimal accuracy degradation by only retraining for 3B tokens. We release our code for training  [^1] Training code: https://github.com/IBM/dolomite-engine and inference  [^2] Inference code: https://github.com/mayank31398/ladder-residual-inference for easier replication of experiments.
Machine Learning, ICML
[…]
Equal contribution
@@ 1 Introduction
With the rapid scaling of Large Language Models (LLMs) (Smith et al., 2022; Workshop et al., 2023; Brown, 2020), the compute and memory requirements for training and inference have grown significantly. Tensor parallelism (TP) (Shoeybi et al., 2020) is a widely used model parallelism technique that partitions the weights and intermediate activations across multiple GPUs. In contrast to pipeline parallelism (Narayanan et al., 2021) and data parallelism (Li et al., 2020), which rely on processing independent batches of user requests on each device, TP enables multiple devices to cooperate to process a single request, therefore in theory allowing infinite scaling given a sufficient number of processors. However, TP requires synchronizing the partitioned intermediate activations across the GPUs. This synchronization is a blocking AllReduce operation on the activations across the GPUs and is therefore bottlenecked by the network communication latency. Even for GPUs connected via fast interconnects (like NVLink (NVIDIA Corporation, 2024)), the communication costs can account for 38% of the latency at inference time when running a 70B transformer with batch size 4 and TP world size of 8.
@@
In current LLMs, communication is blocking because there is a sequential structure between communication and computation in existing model designs: we wait for communication in order to prepare the correct input for the next computation. In the prevalent residual-based architectures, the computation flow can be written as $x_{i+1}=h_{i+1}(x_{i})+x_{i}$, where $x_{i}$ is the residual stream after layer $i$ and $h_{i+1}$ is the computation at layer $i+1$. Notice that the communication of $x_{i}$ needs to be done before executing $h_{i+1}$ if $h_{i}$ is partitioned across devices. Liu et al. (2023b) found that activation changes slowly in Transformer, as the norm of each update $h_{i+1}(x_{i})$ is small compared to the residual. Based on this observation, we hypothesize that maintaining the regular residual connection is enough to restrict the representation shift, and we can feed each module a “stale” input to create overlapping opportunities.
We propose Ladder Residual, a simple change where we reroute the residual stream after module $i-1$ (instead of module $i$) as input to module $i+1$: $x_{i+1}=h_{i+1}(x_{i-1})+x_{i}$. With this design, the computation of $h_{i+1}$ is decoupled from the communication of $x_{i}$, enabling straightforward overlapping to hide the latency of communication. Figure 1 shows how Ladder Residual can be applied to the Transformer architecture. At inference time with TP world size of 8, running a 70B Transformer with Ladder Residual can be around 30% faster than the standard Transformer. In Table 1, we provide the inference speedup on Transformers of different sizes. The proposed Ladder Residual method can also be used to accelerate other forms of parallelism, although we focus on Tensor Parallelism in this paper as it is particularly bottlenecked by the heavy communication. Our method obtains 5-7% training speedup when training an 8B model with 8k context length on 64 H100s with 3D parallelism across the GPUs (Tensor Parallel, Sequence Parallel and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023; Rajbhandari et al., 2020)), but we decide to focus on the inference speed ups since training with pure FSDP is usually faster because weights can be pre-fetched and gradient synchronization using ReduceScatter can be overlapped in FSDP making communications in FSDP non-blocking.
Table 1: Inference speedup from applying Ladder Residual on a Transformer model. The test setup is 1024 prompt length, 512 generated tokens, batch size 4, and TP world size of 8 for 1B, 3B, 8B, 34B, 70B, 176B model and TP world size of 16 for 405B model (using two nodes). The speedup value is calculated by comparing Ladder Transformer with Standard Transformer’s inference throughput in tokens per second. We measure the speedup both with and without P2P communication (modern server nodes usually support P2P communication). Note that disabling P2P communication significantly increases the AllReduce latency.
Model size P2P disabled P2P enabled
@@
Because of the widespread use of Transformer (Vaswani, 2017) based Language models, we focus on applying Ladder Residual on Transformer models in this paper, and we call the resulting model as Ladder Transformer. However, it should be noted that Ladder Residual is compatible with any residual based model architecture. We conduct experiments under two scenarios to verify if we can maintain the same performance as standard Transformer:
@@
  • Post-training adaptation: We take the pretrained Llama-3.1-8B-instruct model (Dubey et al., 2024) and apply Ladder Residual on the upper half layers. We then fine-tune it on 3B tokens to adapt to the representation shift. With this relatively light retraining, we can obtain a hybrid Ladder Llama that is on par with the original Llama model across a variety of multiple-choice and generative tasks.
In our paper, we propose to change the model architecture to accelerate model parallelism without touching low-level kernels, making it easily deployable on any hardware. We show that such architecture modification performs on-par with the standard transformer. As model size grows, multi-gpu or even cross-node serving will become more and more important, and our research provides a fresh perspective on designing a model architecture with parallelism optimizations in mind. Such design can be applied to any architecture that suffers from blocking communication, although in this paper we conduct experiments on Transformer-based language models due to their widespread use and popularity.
2 Tensor Parallelism Background
Tensor parallelism (TP) (Shoeybi et al., 2020) is a widely used technique in distributed training/inference. It partitions weights and activations across devices and performs partial computations on each device. Consider a sequence of 2 linear layers with weight matrices $A$ and $B$ and input activation $X$ that is running on 2 GPUs (TP world size of 2), we split $A$ along the output dimension into $[A_{1},A_{2}]$, and split $B$ along the input dimension into $\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}$. Then the output of the sequence of the 2 linear layers can be computed as $(XA)B=(XA_{1})B_{1}+(XA_{2})B_{2}$ and we effectively partition the computation on the two devices. The final summation requires an AllReduce operation to aggregate the partial sums on each device, which introduces communication overhead. The AllReduce overhead increases with increasing message size and increasing number of devices participating in the AllReduce. A transformer layer consists of an attention block and an MLP block: both can be considered as a sequence of two matrix multiplications and therefore fit into the tensor parallelism paradigm described above. Thus each transformer layer contains 2 AllReduce operations: one for attention and another for MLP. Denoting the input to the $i^{th}$ block as $x_{i-1}$, the transformer can be viewed as the following sequential structure:
@@ 2 Tensor Parallelism Background
where the * denotes a partial-sum that requires an AllReduce to replicate full output across all the GPUs.
A Transformer with N layers needs to perform the AllReduce 2N times and this can account for 38% of the inference latency for a 70B model using TP world size of 8, even with NVLink interconnect. If P2P communication is disabled, AllReduce latency can account for over 50% of the end-to-end latency. Modern nodes connect GPUs via NVLink but usually have no more than 8 GPUs per node, due to limited PCIe lanes or power and cooling constraints in datacenters. There is a steep falloff in communication bandwidth and sharp increase in latency when the communication happens across nodes either over InfiniBand or Ethernet thus making scaling TP practically infeasible outside a node.
Algorithm 1 Ladder Transformer Layer with Tensor Parallelism. Note that the AsyncAllReduce (AAR) returns a handle which is passed to the next layer.
1:function Layer(residual,  attn_out,  mlp_out,                attn_work,  mlp_work)
2:     attn_work.wait()
@@
[…]
15:
16:     returnresidual, attn_out, mlp_out,
17:attn_work, mlp_work
18:end function
3 Ladder Transformer
In this section, we introduce the Ladder Residual architecture when applied to Transformer and benchmark its efficiency under various model sizes and generation setups.
Figure 2: Improvement in end-to-end inference throughput achieved by communication-efficient architectures relative to a standard transformer, benchmarked on Llama-3 70B. Standard refers to the regular Llama-3, and Ladder is Llama-3 with our Ladder Residual architecture. Ladder Residual architecture can achieve up to $29\%$ greater throughput than the standard Transformer. With slower communication (P2P disabled or P2P=0), we observe speedups up to $60\%$. All experiments were conducted on a generation task with $1024$ prompt tokens and $512$ completion tokens. Missing data points indicate CUDA OOM.
@@ 3.1 Architecture description
Note that the residual stream of each block still takes the output from the previous block as usual, this ensures block $i$ can still process information from all previous $i-2$ blocks.
@@
UpperBound-Llama-70B (P2P=1) 30.54 30.00 42.90
Parallel-Llama-70B (P2P=1) 5.42 18.04 21.75
Ladder-Llama-70B (P2P=1) 5.78 23.71 30.79
UpperBound-Llama-70B (P2P=0) 35.84 52.71 110.7
Parallel-Llama-70B (P2P=0) 14.92 28.73 40.07
Ladder-Llama-70B (P2P=0) 6.94 37.71 59.87
@@ 3.2 Inference Implementation
Alignment with Real-World Scenarios: To evaluate the practical benefits of Ladder Residual, we integrated this mechanism into a standard Llama-like Transformer. Building upon gpt-fast (PyTorch Labs, 2024), we partition the weights of the attention and feedforward modules for tensor parallelism to optimize inference speed. We use CUDA graphs (Coleman, 2020) via PyTorch compile (with “reduce-overhead” mode) to generate static computation graphs for both the prefill and decode phase to reduce CPU kernel launch overheads which can be a big bottleneck especially during the decode phase of Transformer inference. PyTorch compile also additionally helps in accelerating inference and reducing the memory footprint for inference by emitting more efficient kernels.
@@ 3.3 Faster Inference with Ladder Residual
In this section, we benchmark Ladder Residual under various scenarios and show that across various model sizes, batch sizes, and TP world sizes, Ladder Transformer can obtain considerable speedup over the standard Transformer. We also benchmark under the case where P2P communication is disabled (for testing in a scenario with increased AllReduce latency), and show that our method can obtain more than 50% speedup. Finally, we consider cross-node Tensor Parallelism, which can be necessary for serving large models like Llama 3.1 405B, and show that Ladder Residual can improve inference throughput by more than $30\%$.
@@ 3.3.1 Setup
  • Parallel Attention and MLP: Following the PaLM parallelization strategy (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021), we fuse the weights of the query, key, value, gate, and up projections into a single matrix. The outputs are then split, and the attention and swiglu are performed in parallel. While not proposed for accelerating Tensor Parallelism, we realize this architecture can effectively cut half of the communication, with the extra benefit of being able to fuse attention and mlp together, therefore we consider it as an alternative to Ladder Residual.
@@
Note that different Transformer-based models usually have slightly different designs. We choose to use Llama-3’s model architecture since it’s one of the most widely used models. The benchmarking results here are a good representative of a variety of different design choices since the communication patterns are mostly the same across various transformer variants.
To simulate various inference scenarios, we select multiple experimental configurations. The prompt length and generation length is fixed to 1024 and 512 respectively, while we vary the tensor parallel world sizes among 1, 2, 4 and 8, and batch sizes among 1, 4, 16 and 64 to understand performance under different generation settings. All benchmarks are done on NVIDIA H100 GPUs.
To evaluate the impact of hardware communication capabilities, we attempt to simulate the performance in the presence of slow interconnects. We set NCCL_P2P_DISABLE=1 to disable the point-to-point communication. This significantly slows down the NCCL communications and allows us to assess the performance of difference algorithms in varying communication environments. We observe a significant slowdown (for 8 GPUs) for AllReduce when disabling P2P communication. In our experiments and figures, we refer to this setting as P2P=0 or P2P disabled. Unless otherwise stated, P2P is always enabled (or P2P=1).